39.21x+2.65=42.03x

can anyone explain to me how to solve this decimal equation and if you can please try to put in 8th grade terms. oh and if you can do you mind solving this one for me too? -7x+32=-21
thankyou!

Found 2 solutions by AnlytcPhil, deepikabansal:
Answer by AnlytcPhil(1810)

39.21x+2.65=42.03x 
can anyone explain to me how to solve this decimal equation and if you can please try to put in 8th grade terms. oh and if you can do you mind solving this one for me too? -7x+32=-21 
thankyou!

The idea is to

1. Get rid of the term which DOES NOT contain x on the LEFT by
   adding its opposite to both sides:
 
39.21x + 2.65 = 42.03x 

Get rid of the + 2.65 by adding -2.65 to both sides

39.21x + 2.65 = 42.03x
       - 2.65 =        - 2.65
-----------------------------
39.21x        = 42.03x - 2.65

Then 

2. Get rid of the term which DOES contain x on the RIGHT by
   adding its opposite to both sides:

       39.21x =  42.03x - 2.65

Get rid of the 42.03x by adding -42.03x to both sides

       39.21x =  42.03x - 2.65
      -42.03x = -42.03x
     --------------------------
       -2.82x =          -2.65

3. Get rid of the coefficient of x, which is -2.82,
   by dividing both sides by -2.82.

      -2.82x     -2.65
     -------- = -------  
      -2.82      -2.82 
 
Cancel the -2.82's on the left

      -2.82x     -2.65
     -------- = -------  
      -2.82      -2.82 

                 -2.65
             x = -------  
                 -2.82 

Divide the numbers on the right

            x = .9397163121

You can round that however you are told, either as
.94, .940, .9397, etc.

----------------------------------------

-7x + 32 = -21


Again the idea is to

1. Get rid of the term which DOES NOT contain x on the 
   LEFT by adding its opposite to both sides:
 
-7x + 32 = -21 

Get rid of the +32 by adding -32 to both sides

-7x + 32 = -21
    - 32 = -32
--------------
-7x      = -53

There are no terms containing x on the right, so this is 
an easier problem and we can skip step 2.

3. Get rid of the coefficient of x, which is -7, by dividing
   both sides by -7.

      -7x     -53
     ----- = -----  
      -7       -7 
 
Cancel the -7's on the left

      -7x     -53
     ----- = -----  
      -7       -7 

             -53
         x = ---  
             -7 

Divide the numbers on the right (cancel the negative signs)

                 53
            x = ----
                  7

You can leave it as 53/7.

Edwin
AnlytcPhil@aol.com

Answer by deepikabansal(5)

Solution: 39.21x+2.65=42.03x
2.65=42.03x-39.21x
2.65=2.82x
x=2.65/2.82
x=0.939
Solution: -7x=-21-32
-7x=-53
x=-53/-7
x=7.571